# Optical pumping of alkali-metal vapor with hyperfine-resolved buffer gas pressure

### 06 Mar 2026 v1

@@ Abstract
Optical pumping is fundamental to high-precision measurement using thermal alkali-metal atoms in vapor cells. In applications such as spin-exchange-relaxation-free magnetometers, buffer gases (e.g., $\text{N}_{2}$ or He) are commonly employed to quench fluorescence and mitigate wall relaxation. In the high-pressure limit (e.g., the $\text{N}_{2}$ pressure $p_{\rm{N}_{2}}>1$ atm), where collisional broadening exceeds the hyperfine splitting of the alkali-metal atoms, optical pumping theory provides a clear description of the angular momentum exchange between photons and atomic spins. However, in many magnetic sensing scenarios, this high-pressure condition is not strictly satisfied, rendering the high-pressure approximation inaccurate. Consequently, a precise quantitative understanding of optical pumping under realistic pressures is critical for determining optimal buffer gas parameters, selecting operating points (e.g., pump frequency and intensity), and enhancing system reliability and stability. To address this, we develop a theory of optical pumping in the quasi-high-pressure regime, where collisional broadening is comparable to the ground-state hyperfine splitting. We demonstrate that optical absorption, spin polarization, and magnetic resonance linewidth in this regime differ significantly from those predicted by the high-pressure limit. Our study extends conventional modeling and offers critical guidance for atomic magnetometry operating under realistic buffer gas pressures.
@@ I Introduction
Depending on the buffer gas pressure, optical pumping operates in distinct regimes, yet a comprehensive theoretical framework covering the intermediate pressures is currently lacking. In the low-pressure (LP) regime with typical buffer gas pressure $\lesssim 10~\mathrm{Torr}$, strong and narrow optical resonances of well resolved hyperfine mutiplets enables refined optical manipulation, such as in atomic clocks based on coherent population trapping (CPT) [24]. In the high-pressure (HP) limit (e.g., buffer gas pressure $1~\mathrm{atm}$), strong collisional quenching and pressure broadening dominate [9, 27, 22]. This treatment renders hyperfine mutiplets unresolved and forms the basis of the spin-temperature distribution (STD) widely used in, e.g., spin-exchange relaxation free (SERF) magnetometers [11, 19]. However, many optical pumping systems including atomic magnetometers operate in the quasi-high-pressure (QHP) regime (typically $\sim 10^{2}~\mathrm{Torr}$). We focus on this intermediate regime, in which the ground-state (GS) hyperfine structure remains spectrally resolvable while the excited-state (ES) structure is unresolved. This leads to spin dynamics that are neither fully captured by the LP models nor the simplified HP approximations. Developing a precise quantitative understanding of the QHP regime is therefore critical. It not only enables the determination of optimal operating points to maximize sensitivity, but also facilitates the assessment of long-term system stability against parameter drifts.
@@
A distinct feature of the QHP regime is the dependence of the absorption cross-section on pumping-light polarization and intensity, which markedly differs from the HP limit. Within our framework, we derive analytical expressions for absorption under both linearly and circularly polarized pumping. We elucidate how the intensity-dependent redistribution of population among hyperfine multiplets reshapes the detuning response of the absorption behavior.
Consequently, the steady-state atomic population in the QHP regime generally deviates from the standard STD. We demonstrate that while the STD can approximate the distribution shape under strong spin-exchange collisions, the effective temperature parameter—namely the spin polarization—requires QHP-specific corrections. Our numerical results regarding spin polarization versus pumping power and frequency reveal that resonant pumping at specific hyperfine levels can yield larger steady-state spin polarization.
Beyond fundamental spin dynamics, we analyze the magneti -resonance linewidth to identify optimal operating conditions for atomic magnetometers. We find that resonant pumping at the lower hyperfine-split levels not only enhances spin polarization but also yields a favorable magnetic-resonance linewidth. This suggests a strategy for optimizing sensitivity through simultaneous improvements in resonance amplitude and linewidth.
The remainder of this paper is organized as follows. Section II defines the QHP regimes based on spectral characteristics. Section III presents the theoretical framework based on the Liouville space master equation. Section IV discusses the unique features of absorption, spin polarization, and magnetic resonance linewidth in the QHP regime. Finally, Section V summarizes the main findings.
@@ II Regimes of buffer gas pressure
Figure 1: Schematic illustration of optical pumping on the alkali-metal D1 line in the HP limit and QHP regime. (a) and (b) show the GS and ES hyperfine multiplets, with different collisional broadening $\Gamma_{\mathrm{brd}}$ illustrated by blue level edges. Each of $F$ and $\bar{F}$ corresponds to either $a=I+1/2$ or b=I1/2b=I-1/2 for nuclear spin $I$. Laser detuning from the D1 line reference frequency is denoted by $\Delta_{0}$, while $\Delta_{F}$ for $F=a$ and $b$ are the detunings corrected from $\Delta_{0}$ by the GS hyperfine splitting. (c) and (d) present the corresponding absorption cross section $\sigma_{0}$ for unpolarized 87Rb, with $\Gamma_{\mathrm{brd}}=3,\,20~\mathrm{GHz}$, respectively. The hyperfine splittings are $\omega^{\mathrm{\{g\}}}_{\mathrm{hf}}=6.8~\mathrm{GHz}$ and $\omega^{\mathrm{\{g\}}}_{\mathrm{hf}}=0.81~\mathrm{GHz}$. Collisional shifts are not included for simplicity. Absorption peaks corresponding to different $F\rightarrow\bar{F}$ transitions are indicated. Doppler broadening $\Gamma_{\mathrm{dop}}=0.57~\mathrm{GHz}$ is included using the Voigt profile, which is negligible in the two regimes.
@@ III.1 Master equation in Liouville space
where $1^{\{\mathrm{g}\}}$ and $1^{\{\mathrm{e}\}}$ are identity operators of the GS and ES subspaces, respectively, and $\omega^{\{\mathrm{eg}\}}$ denotes the reference frequency of the optical transition. In Eqns. (2) & (3), the Hamiltonians $H^{\{\mathrm{g}\}}$ and $H^{\{\mathrm{e}\}}$ are expressed in their eigen-representations with eigen-frequencies $\omega^{\{\mathrm{g}\}}_{F,m_{F}}$ and $\omega^{\{\mathrm{e}\}}_{\bar{F},\bar{m}_{F}}$, together with the eigen-states $|F,m_{F}\rangle$ and $|\bar{F},\bar{m}_{F}\rangle$, respectively, with $F$ ($\bar{F}$) being the total angular momentum quantum number, and $m_{F}$ ($\bar{m}_{F}$) being the magnetic quantum number.
@@
$H_{\mathrm{int}}=V+V^{\dagger}=-\Omega_{\mathrm{I}}\mathbf{d}^{\dagger}\cdot\hat{\mathbf{e}}-\Omega^{*}_{\mathrm{I}}\mathbf{d}\cdot\hat{\mathbf{e}}^{*},$ (4)
with $V$ describing photon absorption and atom excitation with optical Rabi frequency $\Omega_{\mathrm{I}}$, and $V^{\dagger}$ representing the reverse process. Here, $\mathbf{d}$ denotes the dimensionless dipole operators, and $\hat{\mathbf{e}}$ is the unit vector of the positive-frequency classical electric field. For a monochromatic laser beam with frequency $\omega$, the laser detuning with respect to the atomic transition $|F,m_{F}\rangle\leftrightarrow|\bar{F},\bar{m}_{F}\rangle$ is
@@
$\frac{d}{dt}\left|{\rho}^{\{\mathrm{gg}\}}\right)=$ $-\left(iH^{\{\mathrm{g}\}\copyright}+\mathcal{G}^{\{\mathrm{g}\}}_{\mathrm{rel}}+R_{\mathrm{op}}\mathcal{A}_{\mathrm{op}}\right)\left|{\rho}^{\{\mathrm{gg}\}}\right),$ (6)
where $\left|{\rho}^{\{\mathrm{gg}\}}\right)$ is the vectorized GS density matrix, and $X^{\copyright}=X^{\flat}-X^{\dagger\sharp}$ is the commutator superoperator with $X^{\flat}$ and $X^{\sharp}$ the left- and right-translation superoperators [9], corresponding to the operator $X$ in Hilbert space. The GS spin relaxation term $\mathcal{G}^{\{\mathrm{g}\}}_{\mathrm{rel}}$ is
@@
with the spin-destruction rate $\Gamma_{\mathrm{SD}}$, the spin-exchange rate $\Gamma_{\mathrm{ex}}$, and the mean electron spin polarization $(\mathbf{S}\left|{\rho}^{\{\mathrm{gg}\}}\right)=\mathrm{Tr}\left[\mathbf{S}{\rho}^{\{\mathrm{gg}\}}\right]$. The spin-destruction and spin-exchange superoperators $\mathcal{A}_{\mathrm{SD}}$ and $\mathcal{A}_{\mathrm{SE}}$ are
@@
$\mathcal{A}_{\mathrm{rp}}=$ $\Gamma_{\mathrm{q}}\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}\left[{iH^{\{\mathrm{e}\}\copyright}}+\Gamma_{\mathrm{q}}+\Gamma_{\mathrm{JD}}{A}_{\mathrm{JD}}\right]^{-1}\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}},$ (13)
where $\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}$ is the superoperator of quenching process with corresponding quanching rate $\Gamma_{\mathrm{q}}$, and $\mathcal{A}_{\mathrm{JD}}$ is the superoperator of the spin relaxation process of ES with the damping rate $\Gamma_{\mathrm{JD}}$. Analytical expressions of $\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}$ and $\mathcal{A}_{\mathrm{JD}}$ are obtained by the uncoupled irreducible tensor operators, and can be found in Refs. [1, 17]. In Eq. (13), the superoperator $\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}$ is defined as [9]
$\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}=$ $-V^{*}{\otimes}W+W^{*}\otimes V,$ (14)
which describes the optical excitation of the atoms from the GS to the ES. Here, $\otimes$ denotes the Kronecker product between operators in Hilbert space.
@@ III.2.1 Depopulation
The energy denominator matrices $\Delta^{\{\mathrm{eg}\}}+i\Gamma_{\mathrm{brd}}$ in Eqs. (11) and (13) play the central rule in describing optical pumping in different buffer-gas regimes. As discussed in Sec. II, in the HP limit they simply merge into $\Delta_{0}\pm i\Gamma_{\mathrm{brd}}$. Through this approximation, one can obtain $\mathcal{A}^{(\mathrm{HP})}_{\mathrm{dp}}$ [9].
@@
$A^{(\mathrm{HP})}_{\mathrm{dp}}$= $1-\mathbf{s}\cdot\left(\mathbf{S}^{\flat}+\mathbf{S}^{\sharp}\right)-i\mathbf{s}\cdot\mathbf{S}^{\copyright},$ (15)
with $\mathbf{s}=i\hat{\mathbf{e}}\times\hat{\mathbf{e}}^{*}$ the mean photon spin. Here, the identity for D1 line is utilized
$\hat{\mathbf{e}}^{*}\cdot\bm{d}\bm{d}^{\dagger}\cdot\hat{\mathbf{e}}=\frac{1}{2}-\mathbf{s}\cdot\mathbf{S},$ (16)
which expresses how the electric-dipole operator links the GS spin operator.
In the QHP regime of interest, the energy denominators are naturally grouped by $\Delta_{F}\pm i\Gamma_{\mathrm{brd}}$, which characterizes the resolvable $F=a$ and $b$ multiplets and their distinct contributions to optical pumping. In this way, elements of $\delta H^{\{\mathrm{g}\}}$ are simplified as
$\langle F^{\prime}m^{\prime}_{F}|\delta H|Fm_{F}\rangle=\dfrac{\langle F^{\prime}m^{\prime}_{F}|\mathcal{V}^{\dagger}\mathcal{V}|Fm_{F}\rangle}{\Delta_{F}+i\Gamma_{\mathrm{brd}}},$ (17)
with $\Delta_{F}$ being the reference dutuning determined by the F multiplet for $F=a$ or $F=b$
$\Delta_{F}=\Delta_{0}+\omega^{\{\mathrm{g}\}}_{F,0}$. (18)
The amplitude of $\delta H^{\{\mathrm{g}\}}$ is evaluated by $R_{\mathrm{op}}=\sigma_{0}\Phi$, where $\sigma_{0}$ becomes a sum of two Lorentzian line shapes
$\sigma_{0}=A_{0}\left(\frac{[a]}{2[I]}\frac{\Gamma_{\mathrm{brd}}}{\Delta_{{a}}^{2}+\Gamma_{\mathrm{brd}}^{2}}+\frac{[b]}{2[I]}\frac{\Gamma_{\mathrm{brd}}}{\Delta_{{b}}^{2}+\Gamma_{\mathrm{brd}}^{2}}\right).$ (19)
with $[N]=2N+1$ for a quantum number $N$ and $A_{0}$ being the D line constant [12]
$A_{0}=\frac{\left|\Omega_{\mathrm{I}}\right|^{2}}{\Phi}=r_{\mathrm{e}}cf^{\{ge\}}.$ (20)
Here $r_{\mathrm{e}}$ is the classical electron radius, $c$ is the speed of light and $f^{\{ge\}}$ the oscillator strength. Using Eqs. (17) and (19), we obtain the imaginary and real part of $\delta H^{\{\mathrm{g}\}}$ compactly
@@
$\delta H^{\{\mathrm{g}\}}$ $={R_{\mathrm{op}}}\left(\frac{1}{2}-\mathbf{s}\cdot\mathbf{S}\right)\left(2\mathcal{K}_{\mathcal{H}}-i\mathcal{Q}_{\mathcal{H}}\right),$ (21)
which both take the form of $\delta H^{\{\mathrm{g}\}}$ in the HP limit followed by the diagonal matrix $\mathcal{Q}_{\mathcal{H}}$ and $\mathcal{K}_{\mathcal{H}}$
$\mathcal{K}_{\mathcal{H}}$ $=\sum_{F=a,\,b}\sum_{m_{F}}{K}_{F}|Fm_{F}\rangle\langle Fm_{F}|,$ (22)
$\mathcal{Q}_{\mathcal{H}}$ $=\sum_{F=a,\,b}\sum_{m_{F}}{Q}_{F}|Fm_{F}\rangle\langle Fm_{F}|.$ (23)
As shown in Fig. 2, the detuning-dependent factor $Q_{F}$ is
${Q}_{F}$ $=\frac{2[I](\Delta_{F}^{2}+\Gamma_{\mathrm{brd}}^{2})^{-1}}{[a](\Delta_{a}^{2}+\Gamma_{\mathrm{brd}}^{2})^{-1}+[b](\Delta_{b}^{2}+\Gamma_{\mathrm{brd}}^{2})^{-1}}$ (24)
and ${K}_{F}={Q}_{F}\Delta_{F}/\Gamma_{\mathrm{brd}}$. The form in Eq. (21) captures the contribution of each GS hyperfine multiplet to light absorption and optical shift, since the factor $(\Delta_{F}^{2}+\Gamma_{\mathrm{brd}}^{2})^{-1}$ weights the spectral response of the $F$ multiplet.

### 14 Apr 2026 v2

@@ Abstract
Optical pumping is fundamental to high-precision measurement using thermal alkali-metal atoms in vapor cells. In applications such as spin-exchange-relaxation-free magnetometers, buffer gases (e.g., $\text{N}_{2}$ or He) are commonly employed to quench fluorescence and mitigate wall relaxation. In the high-pressure limit (e.g., the $\text{N}_{2}$ pressure $p_{\rm{N}_{2}}>1$ atm), where collisional broadening exceeds the hyperfine splitting of the alkali-metal atoms, optical pumping theory provides a clear description of the angular momentum exchange between photons and atomic spins. However, in many magnetic sensing scenarios, this high-pressure condition is not strictly satisfied, rendering the high-pressure approximation inaccurate. Consequently, a precise quantitative understanding of optical pumping under realistic pressures is critical for determining optimal buffer gas parameters, selecting operating points (e.g., pump frequency and intensity), and enhancing system reliability and stability. To address this, we develop a theory of optical pumping in the quasi-high-pressure regime, where collisional broadening is comparable to the ground-state hyperfine splitting. We demonstrate that optical absorption, spin polarization, and magnetic resonance linewidth in this regime differ significantly from those predicted by the high-pressure limit and offer favorable operating conditions. Our study extends conventional modeling and offers critical guidance for atomic magnetometry operating under realistic buffer gas pressures.
@@ I Introduction
Depending on the buffer gas pressure, optical pumping operates in distinct regimes, yet a comprehensive theoretical framework covering the intermediate pressures is currently lacking. In the low-pressure (LP) regime with typical buffer gas pressure $\lesssim 10~\mathrm{Torr}$, strong and narrow optical resonances of well resolved hyperfine mutiplets enables refined optical manipulation, such as in atomic clocks based on coherent population trapping (CPT) [24]. In the high-pressure (HP) limit (e.g., buffer gas pressure $1~\mathrm{atm}$), strong collisional quenching and pressure broadening dominate [9, 27, 22]. This treatment renders hyperfine mutiplets unresolved and forms the basis of the spin-temperature distribution (STD) widely used in, e.g., spin-exchange relaxation free (SERF) magnetometers [11, 19]. However, many optical pumping systems including atomic magnetometers operate in the quasi-high-pressure (QHP) regime (typically $\sim 10^{2}~\mathrm{Torr}$) [15, 5]. We focus on this intermediate regime, in which the ground-state (GS) hyperfine structure remains spectrally resolvable while the excited-state (ES) structure is unresolved. This leads to spin dynamics that are neither fully captured by the LP models nor the simplified HP approximations. Developing a precise quantitative understanding of the QHP regime is therefore critical. It not only enables the determination of optimal operating points to maximize sensitivity, but also facilitates the assessment of long-term system stability against parameter drifts.
@@
A distinct feature of the QHP regime is the hyperfine-resolved dependence of absorption cross-section on pumping-light polarization and intensity, which markedly differs from the HP limit. Within our framework, we derive analytical expressions for absorption under both linearly and circularly polarized pumping. We elucidate how the redistribution of population among and within hyperfine multiplets reshapes the detuning response of the absorption behavior.
Consequently, driven by the combined redistribution dynamics in the QHP regime, the steady-state atomic population generally deviates from the standard STD. We demonstrate that while the STD can approximate the distribution shape under rapid spin-exchange collisions, the effective temperature parameter—namely the spin polarization—requires QHP-specific corrections. Evaluating these dynamics, we demonstrate that resonant pumping at specific hyperfine levels can yield a larger polarization.
Beyond fundamental spin dynamics, we analyze the magnetic resonance linewidth to identify optimal operating conditions for atomic magnetometers. We find that resonant pumping at the lower hyperfine-split levels not only enhances spin polarization but also yields a favorable magnetic-resonance linewidth. This suggests a strategy for optimizing sensitivity through simultaneous improvements in resonance amplitude and linewidth.
The remainder of this paper is organized as follows. Section II defines the QHP regimes based on spectral characteristics. Section III presents the QHP-specific theoretical framework based on the Liouville space master equation. Section IV discusses the unique features of absorption, spin polarization, and magnetic resonance linewidth in the QHP regime. Finally, Section V summarizes the main findings.
@@ II Regimes of buffer gas pressure
Figure 1: Schematic illustration of optical pumping on the alkali-metal D1 line in the HP limit and QHP regime. (a) and (b) show the GS and ES hyperfine multiplets, with different collisional broadening $\Gamma_{\mathrm{brd}}$ illustrated by blue level edges. Each of $F$ and $\bar{F}$ corresponds to either $a=I+1/2$ or b=I1/2b=I-1/2 for nuclear spin II. Laser detuning from the D1 line reference frequency is denoted by $\Delta_{0}$, while $\Delta_{F}$ for $F=a$ and $b$ are the detunings corrected from $\Delta_{0}$ by the GS hyperfine splitting. (c) and (d) present the corresponding absorption cross section $\sigma_{0}$ for unpolarized 87Rb, with $\Gamma_{\mathrm{brd}}=3,\,20~\mathrm{GHz}$, respectively. The hyperfine splittings are $\omega^{\mathrm{\{g\}}}_{\mathrm{hf}}=6.8~\mathrm{GHz}$ and $\omega^{\mathrm{\{g\}}}_{\mathrm{hf}}=0.81~\mathrm{GHz}$. Collisional shifts are not included for simplicity. Absorption peaks corresponding to different $F\rightarrow\bar{F}$ transitions are indicated. Doppler broadening $\Gamma_{\mathrm{dop}}=0.57~\mathrm{GHz}$ is included using the Voigt profile, which is negligible in the two regimes.
@@ III.1 Master equation in Liouville space
where $1^{\{\mathrm{g}\}}$ and $1^{\{\mathrm{e}\}}$ are identity operators of the GS and ES subspaces, respectively, and $\omega^{\{\mathrm{eg}\}}$ denotes the reference frequency of the optical transition. In Eqs. (2) & (3), the Hamiltonians $H^{\{\mathrm{g}\}}$ and $H^{\{\mathrm{e}\}}$ are expressed in their eigen-representations with eigen-frequencies $\omega^{\{\mathrm{g}\}}_{F,m_{F}}$ and $\omega^{\{\mathrm{e}\}}_{\bar{F},\bar{m}_{F}}$, together with the eigen-states $|F,m_{F}\rangle$ and $|\bar{F},\bar{m}_{F}\rangle$, respectively. In this paper, we assume the magnetic field is so weak that the total angular momentum $F$ and $\bar{F}$ are approximately good quantum numbers, with corresponding magnetic quantum numbers $m_{F}$ and $\bar{m}_{F}$.
@@
$H_{\mathrm{int}}=V+V^{\dagger}=-\Omega_{\mathrm{I}}\bm{d}^{\dagger}\cdot\hat{\bm{e}}-\Omega^{*}_{\mathrm{I}}\bm{d}\cdot\hat{\bm{e}}^{*},$ (4)
with $V$ describing photon absorption and atom excitation with optical Rabi frequency $\Omega_{\mathrm{I}}$, and $V^{\dagger}$ representing the reverse process. Here, $\bm{d}$ denotes the dimensionless dipole operators, and $\hat{\bm{e}}$ is the unit vector of the positive-frequency classical electric field. For a monochromatic laser beam with frequency $\omega$, the laser detuning with respect to the atomic transition $|F,m_{F}\rangle\leftrightarrow|\bar{F},\bar{m}_{F}\rangle$ is
@@
$\frac{d}{dt}\left|{\rho}\right)=$ $-\left(iH^{\{\mathrm{g}\}\copyright}+\mathcal{G}^{\{\mathrm{g}\}}_{\mathrm{rel}}+R_{\mathrm{op}}\mathcal{A}_{\mathrm{op}}\right)\left|\rho\right),$ (6)
where $\left|\rho\right)$ is the vectorized GS density matrix, and $X^{\copyright}=X^{\flat}-X^{\dagger\sharp}$ is the commutator superoperator with $X^{\flat}$ and $X^{\sharp}$ the left- and right-translation superoperators [9], corresponding to the operator $X$ in Hilbert space. The GS spin relaxation term $\mathcal{G}^{\{\mathrm{g}\}}_{\mathrm{rel}}$ is
@@
with the spin-destruction rate $\Gamma_{\mathrm{SD}}$, the spin-exchange rate $\Gamma_{\mathrm{ex}}$, and the expectation value of the electron spin vector $(\mathbf{S}|\rho)=\mathrm{Tr}\left[\mathbf{S}{\rho}\right]$. The spin-destruction and spin-exchange superoperators $\mathcal{A}_{\mathrm{SD}}$ and $\mathcal{A}_{\mathrm{SE}}$ are
@@
$\mathcal{A}_{\mathrm{rp}}=$ $\left(\Gamma_{\mathrm{s}}\mathcal{A}^{\{\mathrm{ge}\}}_{\mathrm{s}}+\Gamma_{\mathrm{q}}\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}\right)$
$\times\left[{iH^{\{\mathrm{e}\}\copyright}}+\Gamma_{\mathrm{s}}+\Gamma_{\mathrm{q}}+\Gamma_{\mathrm{JD}}{A}_{\mathrm{JD}}\right]^{-1}\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}},$ (13)
where $\mathcal{A}_{\mathrm{s}}^{\{\mathrm{ge}\}}$ and $\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}$ are the superoperators of the spontaneous emission and quenching processes with corresponding rates $\Gamma_{\mathrm{s}}$ and $\Gamma_{\mathrm{q}}$. The superaoperator $\mathcal{A}_{\mathrm{JD}}$ describes the spin relaxation process of ES with the damping rate $\Gamma_{\mathrm{JD}}$. Analytical expressions of $\mathcal{A}_{\mathrm{s}}^{\{\mathrm{ge}\}}$ , $\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}$ and $\mathcal{A}_{\mathrm{JD}}$ are derived in Refs. [1, 17]. The superoperator $\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}$ in Eq. (13) is defined as [9]
$R_{\mathrm{op}}\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}=$ $i\left(V^{*}{\otimes}W-W^{*}\otimes V\right),$ (14)
which represents the optical excitation of the atoms from the GS to the ES. Here, $\otimes$ denotes the Kronecker product between operators in Hilbert space.
@@ III.2.1 Depopulation
The energy denominator matrices $\Delta^{\{\mathrm{eg}\}}+i\Gamma_{\mathrm{brd}}$ in Eqs. (12) play a central role in describing optical pumping in different buffer-gas regimes. In the HP limit, the collisional broadening $\Gamma_{\mathrm{brd}}$ is much larger than the difference between the elements of the detuning matrix $\Delta^{\{\mathrm{eg}\}}$, i.e., $\Gamma_{\mathrm{brd}}\gg|\omega_{\bar{F},\bar{m}_{F}}^{\{\mathrm{e}\}}-\omega_{F,m_{F}}^{\{\mathrm{g}\}}|$. In this case, $\Delta^{\{\mathrm{eg}\}}$ is simplified to a single-valued matrix
$\Delta^{\{\mathrm{eg}\}}\approx\Delta_{0}\sum_{\bar{F},\bar{m}_{F},F,m_{F}}|\bar{F},\bar{m}_{F}\rangle\langle F,m_{F}|,$ (15)
and the effective Hamiltonian in Eq. (11) becomes
$\delta H^{\{\mathrm{g}\}}=\frac{V^{\dagger}V}{\Delta_{0}+i\Gamma_{\mathrm{brd}}}.$ (16)
Physically, the approximation in Eq. (15) means that, in the large broadening limit, the different detuning frequencies have a negligible effect on the pumping process, and the alkali-metal atom behaves like a simple two-level atom. With the symmetry property of the electric dipole operator $\hat{\bm{e}}^{*}\cdot\bm{d}\bm{d}^{\dagger}\cdot\hat{\bm{e}}=1/2-K\mathbf{s}\cdot\mathbf{S}$, the HP approximation Eqs. (15) & (16) allows a compact form of the depopulation pumping superoperator [9],
@@
$\mathcal{A}^{(\mathrm{HP})}_{\mathrm{dp}}$= $1-K\mathbf{s}\cdot\left(\mathbf{S}^{\flat}+\mathbf{S}^{\sharp}\right)-\frac{i\Delta_{0}}{\Gamma_{\mathrm{brd}}}K\mathbf{s}\cdot\mathbf{S}^{\copyright},$ (17)
where $K=1$ corresponds to D1 line pumping, with K=1/2K=-1/2 for the D2 line, and $\mathbf{s}=i\hat{\bm{e}}\times\hat{\bm{e}}^{*}$ is the effective photon spin of the pump light. The real and imaginary parts of $\mathcal{A}_{\mathrm{dp}}^{(\mathrm{HP})}$ describe the light absorption and the optical shift, respectively. In Eq. (11), the optical pumping rate $R_{\mathrm{op}}=\bar{\sigma}\Phi$ is proportional to the incident photon flux $\Phi$, and $\bar{\sigma}$ is the intrinsic absorption cross-section of unpolarized atoms (i.e., all the GS levels are equally populated without pumping effect). The intrinsic absorption cross section in the HP limit is a function of detuning $\Delta_{0}$
$\bar{\sigma}_{\mathrm{HP}}=\frac{A}{\pi}\frac{\Gamma_{\mathrm{brd}}}{\Delta_{0}^{2}+\Gamma_{\mathrm{brd}}^{2}}\equiv A\mathcal{L}(\Delta_{0}),$ (18)
where $\mathcal{L}(\Delta)$ is the normalized Lorentzian lineshape function, and $A=\pi r_{\mathrm{e}}cf_{\mathrm{eg}}$ is the area constant with the classical electron radius $r_{\mathrm{e}}$, the speed of light $c$ and the atom oscillator strength $f^{\{ge\}}$.
In the QHP regime of interest, the ES hyperfine splittings and all the Larmor frequencies in weak magnetic fields are neglected, i.e., $\omega_{\bar{F},\bar{m}_{F}}^{\{\mathrm{e}\}}\approx 0$ and $\omega_{F,m_{F}}^{\{\mathrm{g}\}}\approx\omega_{F,0}^{\{\mathrm{g}\}}$, but the GS hyperfine splitting is kept. In this case, the elements of the detuning matrix $\Delta^{\{\mathrm{eg}\}}$ are grouped by two values $\Delta_{F}$ for $F=a$ or $b$, i.e.
$\Delta^{\{\mathrm{eg}\}}\approx\sum_{\bar{F},\bar{m}_{F},F,m_{F}}\Delta_{F}|\bar{F}\bar{m}_{F}\rangle\langle F,m_{F}|$ (19)
where $\Delta_{F}=\Delta_{0}+\omega^{\{\mathrm{g}\}}_{F,0}$ is the frequency detuning with respect to hyperfine split levels. With this coarse-grained modeling of the detuning matrix, the effective Hamiltonian $\delta H^{\{\mathrm{g}\}}$ becomes
$\delta H^{\{\mathrm{g}\}}=$ $\sum_{F=a,b}\delta H^{\{\mathrm{g}\}}_{F}\Pi_{F},$ (20)
$\delta H^{\{\mathrm{g}\}}_{F}$= $\frac{V^{\dagger}V}{\Delta_{F}+i\Gamma_{\mathrm{brd}}},$ (21)
where $\Pi_{F}=\sum_{m_{F}}|F,m_{F}\rangle\langle F,m_{F}|$ is the projection operator of the GS subspace with a given total angular momentum quantum number $F$, and $\delta H_{F}^{\{\mathrm{g}\}}$ is the $F$-resolved effective Hamiltonian.
With the QHP correction in Eq. (19), we show that the depopulation superoperator $\mathcal{A}_{\mathrm{dp}}$ becomes
@@
$\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}=\sum_{F}\left[1-K\mathbf{s}\cdot(\mathbf{S}^{\flat}+\mathbf{S}^{\sharp})-\frac{i\Delta_{F}}{\Gamma_{\mathrm{brd}}}K\mathbf{s}\cdot\mathbf{S}^{\copyright}\right]Q_{F}\mathcal{P}_{F}.$ (22)
See Appendix A for detailed proof. In comparison to the HP case Eq. (17), the hyperfine-resolved GS brings about the following corrections: (i) the overall detuning $\Delta_{0}$ is replaced by the $F$-resolved detuning $\Delta_{F}$; (ii) a weighting factor $Q_{F}$ is introduced as
$Q_{F}(\Delta_{F})=\frac{\mathcal{L}(\Delta_{F})}{x_{a}\mathcal{L}(\Delta_{a})+x_{b}\mathcal{L}(\Delta_{b})},$ (23)
where $x_{F}=(2F+1)/(2(2I+1))$ is the degeneracy weighting factor; and (iii) the expression in Eq. (22) is subjected in the Zeeman subspace with the projection operator in the Liouville space
$\mathcal{P}_{F}=$ $\sum_{m_{F},m^{\prime}_{F}}|F,m_{F};F,m^{\prime}_{F})(F,m_{F};F,m^{\prime}_{F}|,$ (24)
$\mathcal{P}_{\mathrm{Z}}=$ $\mathcal{P}_{a}+\mathcal{P}_{b}.$ (25)
The intrinsic absorption cross-section in the QHP regime should be modified as
$\bar{\sigma}_{\mathrm{QHP}}=A\left[x_{a}\mathcal{L}(\Delta_{a})+x_{b}\mathcal{L}(\Delta_{b})\right].$ (26)